OpenStudy (anonymous):

the range of y=arcsinx is(-π/2, π/2) . A. True B. False

OpenStudy (anonymous):

true. really it is true

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

oh wait. actually it is \[[-\frac{\pi}{2},\frac{\pi}{2}]\]

OpenStudy (anonymous):

closed brackets not open ones

OpenStudy (anonymous):

how is kassia's pi different then satellite's

OpenStudy (anonymous):

so if you have open brackets ( ) instead of [ ] it is false

OpenStudy (anonymous):

i better go

OpenStudy (anonymous):

kassia what do you type for pi